OpenStudy (anonymous):

Use substitution to solve the system of equations. 2x-6y=8 -4x+12y=-16 A. (0, -4/3) B. (4/3,0) C. (4,0) D. infinitely many solutions

OpenStudy (igreen):

Substitution goes like this: Rearrange the first equation so that it equals \(x\). 2x - 6y = 8 Add 6y to both sides: 2x = 6y + 8 Divide 2 to both sides: x = 3y + 4

OpenStudy (igreen):

Since \(x\) is \(3y + 4\), we can input \(3y + 4\) in the second equation.

OpenStudy (igreen):

\(-4x+12y=-16\) \(-4(3y + 4)+12y=-16\) Solve for \(y\): \(-12y - 16 + 12y = -16\) Add 16 to both sides: \(-12y + 12y = 0\) \(0 = 0\) Well, from here you can already tell that it has infinitely many solutions :)

OpenStudy (igreen):

Welcome to Open Study! You can give medals by clicking 'Best Response'. @Brandi0101 \(\bbox[5pt, #000000 ,border:5px solid #00FF00]{\bf\huge \color{#00FF00}\sim \! \! \! \! \color{#00FF00}i\it\Huge\color{#00FF00}G\color{#00FF00}r\color{#00FF00}e\color{#00FF00}e\color{#00FF00}n\bf\color{#00FF00}?}\)